# How Does Perfect Fitting Affect Representation Learning? On the Training Dynamics of Representations in Deep Neural Networks

### v1

@@ 3.3 Analysis of the CKA Results
Figure 3: CKA evaluations for ViT-B/16, trained on CIFAR-10 without label noise.
@@
Representations at the first layer are more similar to random representations than to good early-stopping representations.
@@
The CKA diagrams of the first layer in ResNet-18 trained on CIFAR-10 without/with label noise (Figs. 1a, 2a) show higher CKA values than the other layers, suggesting that representations of the first layer do not change much compared to their random initialization. Yet, within this range of high CKA values, the diagrams show that in the perfect fitting regime of training the first layer representation is more similar to the random representations at the initialization and early in training rather than to representations in the beneficial early-stopping range of training (i.e., at epochs where the test error is relatively low and perfect fitting did not start yet, according to Figs. 1g, 2g). Note that the similarity to the random representations is not only to the specific initialization of the training process, but also to other random representations unrelated to the specific training process; we examine this further in Appendix H.1 and Fig. 19. While most of our results are for the Adam optimizer, for the SGD optimizer (Figs. 13a, 14a, 15a) the first layer representations behave somewhat differently, although they still have very high CKA similarity during the entire training process.
Importantly, the representations of the first layer have increased similarity to the random initialization at the perfect fitting regime only if the DNN is wide enough. For example, see Figs. 8a, 9a where we examine ResNet-18 of smaller widths (width parameters $k=32,42$) that still achieve perfect fitting (and show epoch-wise double descent) but do not have increased similarity to initialization at the end of training. When the DNN is wider (width parameter $k=52$), in Fig. 10a, the behavior is more similar to our results in Fig. 2a for the standard width of ResNet-18 (width parameter $k=64$).
Training the standard-width ResNet-18 using weight decay regularization with the Adam optimizer (Fig. 16a) shows similar behavior of the first layer representations as for smaller width ResNet-18 trained without weight decay (e.g., Fig. 10a). This can be explained by the effective reduction of the model complexity due to the weight decay regularization, making it similar to the model complexity of unregularized smaller model. Note that the weight decay regularization lowers the CKA similarity of the first layer representations (Fig. 16a) in a few percentages compared to training without weight decay (even of smaller width ResNet-18 models); this suggests that weight decay can increase to some extent the evolution of the first layer representations.
Deeper layers evolve much more when epoch-wise double descent occurs.
Figures 1, 2 consider the same learning setting for ResNet-18, except for the single difference of having noiseless versus 20% label noise in the CIFAR-10 dataset. Accordingly, an epoch-wise double descent in the test error occurs only for the noisy case (compare Figs. 1g, 2g). Visually comparing the CKA diagrams of the two cases (Figs. 1,2) shows that the representations at the deeper layers evolve much more in the training when an epoch-wise double descent occurs. This is visually emphasized by the darker color (i.e., lower similarity) in the heatmaps of the fourth ResNet Block (Figs. 1e, 2e) and the last FC layer (Figs. 1f, 2f), reflecting lower CKA similarity among nearby training epochs due to excessive instability and/or evolution of the representations of these layers.
Representations can have one or two transition points in their training evolution, creating “similarity blocks” along the main diagonal of the similarity diagram.
For ResNet, this behavior is especially visible in settings with label noise and epoch-wise double descent (Fig. 2). For example, see Figs. 2c-2f where at each layer “similarity blocks” of relatively lighter color appear around the main diagonal of the diagram. Importantly note that different ResNet layers can differ in the number of “similarity blocks” as well as the range of epochs in which the “similarity blocks” emerge. Specifically, the earlier ResNet layers (Figs. 2b-2d) have larger “similarity blocks” that do not seem to be affected by the perfect fitting regime of training; in contrast, the deeper layers (Figs. 2e-2f) have smaller (or medium size) “similarity blocks” that seem to be more affected by the perfect fitting regime (in particular note that the last FC layer has a small “similarity block” towards the end of training which is very dissimilar to the entire training process; see this also for ResNet-50 in Fig. 12f). In the noiseless setting (where there is no epoch-wise double descent) of Fig. 1, the “similarity blocks” appear more weakly than in the noisy case; this emphasizes the more segmented nature of representation learning in ResNet when the epoch-wise double descent occurs. Yet, the in the noiseless case of ResNet-18, there are “similarity blocks” that are aligned with the perfect fitting regime (i.e., they start near to epoch 1000 where the perfect fitting starts, see Figs. 2e, 2f). Moreover, interestingly, training ResNet-18 with SGD (Figs. 13, 14, 15) has clearer “similarity blocks” and transition points between them than in training with Adam.
ViT has clear “similarity blocks” and transition points which are aligned with the perfect fitting regime.
The phenomenon of “similarity blocks” is somewhat visibly clearer in ViT than in ResNet. Specifically, in ViT training the “similarity blocks” are well observed for both noiseless (Fig. 3) and noisy (Fig. 17) datasets. Specifically, the ViT encoder blocks at various depths have “similarity blocks” which are aligned with the perfect fitting regime (i.e., they start around epoch 1000 where perfect fitting also starts; see Figs. 3e, 3f, 17j, 17m, noting that at the deep layers these “similarity blocks” are visibly clearer).
@@
4 Decision Regions Similarity of Linear Classifier Probes of Layers During Training
@@
Following the insights we obtained using the CKA evaluations, we now turn to a second, significantly different, approach to analyze the training dynamics of the learned representations and features in the DNN layers. Specifically, in this section, we propose a new analysis approach that can be of independent interest also beyond the research questions of this paper.
@@ 4.1 Training Linear Classifier Probes for Layers Throughout the Training Process
From the complete training process, we define a relatively dense subset of training epochs $\mathcal{T}\subset\{0,1,\dots,3999\}$ for which we examine various layers in the trained DNN. As the notations earlier in this paper, for layer $L$ at epoch $t\in\mathcal{T}$, the mapping from the input to the layer’s representation is $f_{L}^{(t)}:\mathcal{X}\rightarrow\mathbb{R}^{p_{L}}$. We use $f_{L}^{(t)}$ (i.e., the DNN segment from the input to layer $L$) as a fixed feature map, and add a trainable linear layer on top of it to act as a probe. Namely, we train the weights $\mathbf{W}\in\mathbb{R}^{|\mathcal{Y}|\times p_{L}}$ and bias $\mathbf{b}\in\mathbb{R}^{|\mathcal{Y}|}$ for the linear classifier probe
$\widetilde{{\sf Probe}}_{L}^{(t)}\left({f_{L}^{(t)}\left(\mathbf{x}\right)}\right)={\sf softmax}\left({\mathbf{W}}f_{L}^{(t)}\left(\mathbf{x}\right)+\mathbf{b}\right)$ (3)
@@
where $\mathcal{Y}$ is the label domain of the original classification task. The training of the linear probe is completely separate from the training of the original DNN that $f_{L}^{(t)}$ is taken from. Moreover, linear probes at different training epochs (of the original DNN) are trained separately and do not share their weights $\mathbf{W}$ and bias $\mathbf{b}$. See more details on the probe training in Appendix F.3. Note that the linear probe layer in (4) maps the $p_{L}$-dimensional representation of $f_{L}^{(t)}$ to a score vector of classes (these classes are as in the classification task of the original DNN, and the probe’s score vector also goes through a softmax function). The predicted class label from $\mathcal{Y}$ is the one with the maximal softmax-based class score, namely, by denoting the $j^{\sf th}$ class score as $\widetilde{{\sf Probe}}_{L}^{(t)}\left({f_{L}^{(t)}\left(\mathbf{x}\right)}\right)[j]$ we have
@@
${\sf Probe}_{L}^{(t)}\left({f_{L}^{(t)}\left(\mathbf{x}\right)}\right)={\sf argmax}_{j\in\{1,\dots,|\mathcal{Y}|\}}\widetilde{{\sf Probe}}_{L}^{(t)}\left({f_{L}^{(t)}\left(\mathbf{x}\right)}\right)[j].$ (4)
4.2 Measuring Similarity of Decision Regions of Linear Classifier Probes
Figure 4: Decision regions similarity of linear classifier probes. Evaluations for ResNet-18 (of standard width), trained on CIFAR-10 without label noise. Each subfigure shows the DRS of a specific layer in the ResNet-18 during its training. The error curves are provided in Fig. 1g.
Figure 5: Decision regions similarity (DRS) of linear classifier probes. Evaluations for ResNet-18 (of standard width), trained on CIFAR-10 with 20% label noise. Each subfigure shows the DRS of a specific layer in the ResNet-18 during its training. The test and error curves are provided in Fig. 2g.
@@ 4.2 Measuring Similarity of Decision Regions of Linear Classifier Probes
Linear classifier probes are often used for analyzing the quality of the learned features by evaluating their usefulness for the original [3, 11] or another classification task [10]. In this paper, we are interested in understanding the evolution of representations during training; accordingly, we propose here a new probing-based similarity measure that reflects the overall representation. Specifically, we propose to compare representations of the same layer at two different training epochs by evaluating the similarity of the decision regions of the linear classifier probes of the layer at the respective training epochs.
@@
Evaluating similarity of decision regions was proposed in [12] for comparing fully-trained DNNs in their standard classification functionality of the entire DNN (i.e., not its layers and without probes). Here, we extend this decision-region similarity evaluation to the completely different analysis task of comparing linear probes that correspond to intermediate layers during their training.
@@
Given two linear classifier probes of the same layer $L$ at different training epochs $t,t^{\prime}$, we numerically estimate the relative portion of the input space for which their predictions agree. As in [12], we define $N_{Q}=500$ planes that reside in the input space, each plane is defined by randomly choosing three training examples and computing the plane that they span. See more details in Appendix F.3. The $j^{\sf th}$ plane, $j\in\{1,\dots,N_{Q}\}$, is denoted as $\mathcal{Q}_{j}\subset\mathcal{X}$. $\mathcal{Q}_{j}^{\sf discrete}$ is a discrete, finite version of $\mathcal{Q}_{j}$, which is defined for the numerical evaluation based on a uniformly-spaced 2D grid of $|\mathcal{Q}_{j}^{\sf discrete}|=2500$ points. Each $\mathbf{x}\in\mathcal{Q}_{j}^{\sf discrete}$ is an image, sampled from the plane, that can be given to $f_{L}^{(t)}$ and $f_{L}^{(t^{\prime})}$ as input. Then, the similarity of the decision regions of the probes of layer $L$ at epochs $t,t^{\prime}$ is
@@
${\sf DRS}\left(f_{L}^{(t)},f_{L}^{(t^{\prime})}\right)\triangleq\frac{1}{N_{Q}}\sum_{j=1}^{N_{Q}}\sum_{\mathbf{x}\in\mathcal{Q}_{j}^{\sf discrete}}\frac{\mathbb{I}\left({\sf Probe}_{L}^{(t)}\left({f_{L}^{(t)}\left(\mathbf{x}\right)}\right)={\sf Probe}_{L}^{(t^{\prime})}\left({f_{L}^{(t^{\prime})}\left(\mathbf{x}\right)}\right)\right)}{|\mathcal{Q}_{j}^{\sf discrete}|}$ (5)
@@
where $\mathbb{I}({\sf condition})$ is an indicator function that returns 1 if the ${\sf condition}$ is satisfied, and 0 otherwise. According to (5), the value of ${\sf DRS}\left(f_{L}^{(t)},f_{L}^{(t^{\prime})}\right)$ is between 0 and 1, where a higher value reflects higher similarity and agreement of the decision regions of the linear probes.
@@
The foundational idea that we propose here is that similar representations lead to similar decision regions of the respective linear classifier probes. In this approach, the intricacy of comparing the representations of different models (from different training epochs) is first addressed by the training of linear classifier probes, which interpret the representational structure to achieve their own training objectives. Clearly, the functionality that the probes learn reflect and depend on the representational structure of their inputs, while having the good merits (for representational similarity evaluation) of insensitivity to the particular order of neurons, orthogonal transformations, or uniform scaling (it should be noted though that there is some level of inherent randomness in the training process of the probes). Comparing the decision regions of the probes is a way to compare their functionality and, by that, to compare the representational structure they get as their inputs. Clearly, this is an inherently different approach to measure representational similarity than using CKA as in Section 3; indeed, we will now see it can shed additional light on the evolution of representations in training and further support insights we obtained using the CKA.
4.3 Analysis of the Linear Classifier Probes Results

### v2

@@ 3.3 Analysis of the CKA Results
Figure 2: CKA evaluations for ViT-B/16, trained using Adam on CIFAR-10 without label noise.
Representations can have one or two transition points in their training evolution, creating similarity blocks of relatively lighter color around the main diagonal of the similarity diagram. Different layers can differ in the number of similarity blocks as well as the range of epochs in which the similarity blocks emerge. We observe that similarity blocks may be aligned with the training phase transitions in some of the DNN layers, but not in all of them; this gives a new perspective on the gradient-based analysis by Maini et al. (2023) that memorization can be due to some (possibly arbitrary) layers or neurons in the DNN. For example, the similarity blocks are aligned with the training phases in the first conv layer and the fourth ResNet Block output for ResNet-18 training (Figs. 1a, 1e); and in the outputs of Encoder Blocks 4-12 in ViT training (see Figs. 2, D.12).
Our results show that similarity blocks can be more pronounced (i.e., reflecting relatively-quick change that causes higher dissimilarity of representations before and after a time point in training) depending on the DNN architecture and optimizer – we will now explain this in more detail.
3.3.1 ResNet vs. ViT
Training phase transitions are much more abrupt and synchronized across layers, and similarity blocks are more clearly identified, in ViT than in ResNet.
The similarity blocks are usually visibly clearer in ViT than in ResNet. Specifically, in ViT training the similarity blocks are well observed for both noiseless (Fig. 2, D.20) and noisy (Fig. D.12) datasets. Remarkably, our results show that the training dynamics of representations is highly synchronized across the ViT layers; this is due to the strong effect of skip connections in ViT, as demonstrated by Raghu et al. (2021) for representational similarity of fully trained ViTs – we, therefore, significantly extend their finding by showing that the entire representation learning process is synchronized. These synchronized dynamics are also aligned with the training phases, implying that memorization of atypical/noisy examples and perfect fitting can more clearly explain the representation learning evolution in ViT than in ResNet.
3.3.2 Adam vs. SGD
Training phase transitions are much more abrupt, and similarity blocks are more clearly identified, in SGD training than in Adam training.
Training ResNet-18 with SGD (e.g., Figs. D.9, D.10, D.15) has clearer similarity blocks and transition points between them than in training with Adam (e.g., Figs. 1, D.1). This implies that memorization of atypical/noisy examples has a greater effect on representation learning dynamics in SGD than in Adam training. In Appendix D.4 we directly compare Adam and SGD training processes using CKA and discuss the results.
3.3.3 First Layer in Convolutional Networks
@@
In the perfect fitting regime of training with Adam, representations at the first layer are more similar to random representations than to good early-stopping representations.
@@
The CKA diagrams of the first layer in ResNet-18 trained using Adam without/with label noise (see Figs. 1a, D.1a, D.2a for CIFAR-10; Figs. D.18a for SVHN) show higher CKA values than the other layers, suggesting that representations of the first layer do not change much compared to their random initialization. Yet, within this range of high CKA values, the diagrams show that in the perfect fitting regime of training the first layer representation is more similar to the random representations at the initialization and early in training rather than to representations in the beneficial early-stopping range of training (i.e., at epochs where the test error is relatively low and perfect fitting did not start yet).
Note that the similarity to the random representations is not only to the specific initialization of the training process, but also to other random representations unrelated to the specific training process; we examine this further in Appendix D.1 and Fig. D.13.
For the SGD optimizer (e.g., Figs. D.8a, D.9a, D.10a) the first layer representations behave somewhat differently than for Adam, although they still have very high CKA similarity during the entire training process. The lack of this phenomenon in SGD training suggests that its existence may require adaptive normalization of gradients such as implemented in Adam (and SGD does not have).
It was shown recently by Cattaneo et al. (2024) that, in contrast to SGD, the implicit regularization of Adam can significantly change throughout the training – e.g., implicit anti-regularization that late in training turns into regularization. Accordingly, we believe that our observations here on the first layer representational behavior in Adam training can be potentially due to changes in implicit regularization. Specifically, as discussed by Cattaneo et al. (2024), when the componentwise gradient magnitudes become lower than the Adam’s numerical stability parameter $\epsilon$, the normalizing adaptivity of gradients becomes less effective. Indeed, we show in Figs. D.21f that our first-layer phenomenon does not exist in Adam training with a large $\epsilon$ ($0.01$ instead of the standard $10^{-8}$) that effectively disables the adaptive gradient normalization (see gradient histogram in Fig. D.21b); this further shows the implicit role that the adaptive gradient normalization has on representation learning dynamics. Importantly, Cattaneo et al. (2024) focus on mathematical analysis of the learned model parameters and their gradients (and not representational similarity) and their empirical results for DNNs use full batch Adam; in contrast, we focus here on layerwise representational similarity in realistic settings of mini-batch Adam, therefore, our findings here consider more intricate DNN training procedures and significantly add to those by Cattaneo et al. (2024).
The representations of the first layer have increased similarity to the random initialization at the perfect fitting regime only if the DNN is wide enough (Figs. D.3a, D.4a); we discuss this in more detail in Appendix D.2. The phenomenon is observed not only for ResNet-18, but also for a two-layer convolutional model (trained on 1k examples from Fashion MNIST, Fig. D.17). Accordingly, we believe this phenomenon may occur in Adam training of overparameterized (sufficiently wide) convolutional networks.
@@
4 Task-Specific Representations: Decision Regions Similarity of Linear Classifier Probes of Layers During Training
Figure 3: Decision regions similarity of linear classifier probes. Evaluations for ResNet-18 trained on CIFAR-10 without label noise. Each subfigure shows the DRS of a specific layer in the ResNet-18 during its training. The error curves are provided in Fig. 1g.
Figure 4: Visualization of decision regions of the linear probe of the first convolution layer in ResNet-18 (trained on CIFAR-10).
@@
Following our CKA-based insights, we now turn to a significantly different approach to analyze training dynamics of learned representations in DNN layers. In this section, we propose a new analysis approach that elucidates the task-specific similarity among representations during training.
@@ 4.1 Training Linear Classifier Probes for Layers Throughout the Training Process
From the complete training process, we define a relatively dense subset of training epochs $\mathcal{T}\subset\{0,1,\dots,3999\}$ for which we examine various layers in the trained DNN. We denote again that for layer $L$ at epoch $t\in\mathcal{T}$, the mapping from the input to the layer’s representation is $f_{L}^{(t)}:\mathcal{X}\rightarrow\mathbb{R}^{p_{L}}$. We use $f_{L}^{(t)}$ (i.e., the DNN segment from the input to layer $L$) as a fixed feature map, and add a trainable linear layer on top of it to act as a probe. Namely, we train the weights $\mathbf{W}\in\mathbb{R}^{|\mathcal{Y}|\times p_{L}}$ for the linear classifier probe
$\widetilde{{\sf Probe}}_{L}^{(t)}\left({f_{L}^{(t)}\left(\mathbf{x}\right)}\right)={\sf softmax}\left({\mathbf{W}}f_{L}^{(t)}\left(\mathbf{x}\right)\right)$ (2)
@@
where $\mathcal{Y}$ is the label domain of the original classification task. The training of the linear probe is completely separate from the training of the original DNN that $f_{L}^{(t)}$ is taken from. Moreover, linear probes at different training epochs (of the original DNN) are trained separately and do not share their weights $\mathbf{W}$. See more details on the probe training in Appendix B.5. Note that the linear probe layer in (3) maps the $p_{L}$-dimensional representation of $f_{L}^{(t)}$ to a score vector of classes (these classes are as in the classification task of the original DNN, and the probe’s score vector also goes through a softmax function). The predicted class label from $\mathcal{Y}$ is the one with the maximal softmax-based class score, namely, by denoting the $j^{\sf th}$ class score as $\widetilde{{\sf Probe}}_{L}^{(t)}\left({f_{L}^{(t)}\left(\mathbf{x}\right)}\right)[j]$ we have
@@
${\sf Probe}_{L}^{(t)}\left({f_{L}^{(t)}\left(\mathbf{x}\right)}\right)=\operatorname*{argmax}_{j\in\{1,\dots,|\mathcal{Y}|\}}\widetilde{{\sf Probe}}_{L}^{(t)}\left({f_{L}^{(t)}\left(\mathbf{x}\right)}\right)[j].$ (3)
4.2 Measuring Similarity of Probe Decision Regions
Figure 5: Visualization of decision regions of the layerwise linear probes and full DNN prediction in ViT-B/16 (trained on CIFAR-10).
@@ 4.2 Measuring Similarity of Probe Decision Regions
Linear classifier probes are often used for analyzing the quality of the learned features by evaluating their usefulness for the original (Alain & Bengio, 2017; Raghu et al., 2021) or another classification task (Hermann & Lampinen, 2020). In this paper, we are interested in understanding the evolution of representations during training; accordingly, we propose here a new probing-based similarity measure that reflects the overall representation. Specifically, we propose to compare representations of the same layer at two different training epochs by evaluating the similarity of the decision regions of the linear classifier probes of the layer at the respective training epochs.
This Decision Region Similarity (DRS) approach considers the representations in a task-specific manner, as the probes train to classify the layer representations to the same set of classes the original DNN has. The DRS approach is importantly different from the CKA that does not explicitly use the specific task in the representation analysis.
@@
Evaluating similarity of decision regions was proposed by Somepalli et al. (2022) for comparing fully-trained DNNs in their standard classification functionality of the DNN prediction output (i.e., not its layers and without probes). Here, we extend this decision-region similarity evaluation to the completely different analysis task of comparing linear probes of intermediate layers during their training.
4.2.1 The DRS Computation Procedure
@@
Given two linear classifier probes of the same layer $L$ at different training epochs $t,t^{\prime}$, we numerically estimate the relative portion of the input space for which their predictions agree. As Somepalli et al. (2022), we define $N_{Q}=500$ planes that reside in the input space, each plane is defined by randomly choosing three training examples and computing the plane that they span. See more details in Appendix B.5. The $j^{\sf th}$ plane, $j\in\{1,\dots,N_{Q}\}$, is denoted as $\mathcal{Q}_{j}\subset\mathcal{X}$. $\mathcal{Q}_{j}^{\sf discrete}$ is a discrete, finite version of $\mathcal{Q}_{j}$, which is defined for the numerical evaluation based on a uniformly-spaced 2D grid of $|\mathcal{Q}_{j}^{\sf discrete}|=2500$ points. Each $\mathbf{x}\in\mathcal{Q}_{j}^{\sf discrete}$ is an image, sampled from the plane, that can be given to $f_{L}^{(t)}$ and $f_{L}^{(t^{\prime})}$ as input. Then, the similarity of the decision regions of the probes of layer $L$ at epochs $t,t^{\prime}$ is
${\scriptstyle{\sf DRS}\left(f_{L}^{(t)},f_{L}^{(t^{\prime})}\right)\triangleq}$ (4)
@@
${\scriptstyle\frac{\sum_{j=1}^{N_{Q}}\sum_{\mathbf{x}\in\mathcal{Q}_{j}^{\sf discrete}}\mathbb{I}\left({\sf Probe}_{L}^{(t)}\left({f_{L}^{(t)}\left(\mathbf{x}\right)}\right)={\sf Probe}_{L}^{(t^{\prime})}\left({f_{L}^{(t^{\prime})}\left(\mathbf{x}\right)}\right)\right)}{N_{Q}\cdot|\mathcal{Q}_{j}^{\sf discrete}|}}$
@@
where $\mathbb{I}({\sf condition})$ is an indicator function that returns 1 if the ${\sf condition}$ is satisfied, and 0 otherwise. The ${\sf DRS}\left(f_{L}^{(t)},f_{L}^{(t^{\prime})}\right)$ value is between 0 and 1, where a higher value reflects higher similarity and agreement of the decision regions of the linear probes.
@@
Our proposed idea here is that similar representations lead to similar decision regions of the respective linear classifier probes. In this approach, the intricacy of comparing the representations of different models (from different training epochs) is first addressed by the training of linear classifier probes, which interpret the representational structure to achieve their own training objectives. Clearly, the functionality that the probes learn reflects and depends on the representational structure of their inputs. It should be noted though that there is some level of inherent randomness in the training process of the probes. Comparing the decision regions of the probes is a way to compare their functionality and, by that, to compare the task-specific representational structure they get as their inputs. Importantly, the DRS is an interpretable approach, as the decision regions can be visualized and and further elucidate the representation learning dynamics beyond the CKA capabilities.
4.2.2 Analysis of the DRS Results